determine the EOQ when the annual demand is 250000 the carrying cost her unit is 30 and ordering cost 750.if the warehouse constrain can only hold 7500 units.if the order is constrained by this what is the annual cost
Answer #1

EOQ = Square root ((2xAnnual demand in Units x ordering cost) /Carrying cost per unit)

= Square root ((2x250000x750)/30)

= Square root (12,500,000)

= 3536 units

Number of order to be placed in a year = 250,000/3536 = 70.72 ~71 orders

Average inventory = 3536/2 = 1768

Annual Carrying cost =(1768*30) =53,040

Total ordering cost =71*750 =53,250

Total annual cost= $106,290